The Advanced Photon Source
a U.S. Department of Energy Office of Science User Facility

APS-UES Meeting Minutes - 10/24/2016

Attendees: Bruce Glagola, Tom Barkalow, Clay White, Bruno Fieramosca, Wendy VanWingeren, Dean Wyncott, and Patti Pedergnana

Bruce Glagola

  • The laboratory placard information should be in.  Shane will get the final list to Tom soon.
  • Our group no longer has to attend Maintenance Monday meetings unless directly contacted.

Tom Barkalow

  • There is a Work Planning and Control Assessment going on this week and next to prepare for the DOE Assessment coming the first week of April next year.  Bruce will talk to the assessors about the experiments running the week of November 2nd.
  • A management assessment of the experiment hazard classes is needed before next April.
  • There were independent reviews done during August and from that came recommendations on better communication and corrective action plans are being developed.  We may see action items coming our way.

Patti Pedergnana

  • The MOM group is at 11-ID and 16-ID today to repair air leaks in a few door pneumatic boxes.
  • 12-ID-D would like the 433/434 inner truck lock door LOTO procedure and door controls locked out for temperature sensitive experiments.

Dean Wyncott

  • 35-ID had a BSDRSC for the installation of the laser transport and mirror.  This will be installed over the next couple weeks and the beamline is hoping for a shielding verification on November 21st.
  • 34-ID’s undulator goes into System Manager Mode with a system error from time to time over the last couple years.  This was reported to John Grimmer who is aware of the reasons this error occurs.
  • Dean met with Zhan Zhang about installing a LN2 deck on the top of the 33-ID-A station.  They will have it built by Central Shops if they are able to do the work.
  • 35-ID-E has a safety board on a door near its control area on the inboard side of the station.  Please do not post the EHCP on the outboard, experimental hall side, door.
  • Dean submitted his BEQ and had questions about chemicals in CORAL.

Wendy VanWingeren

  • 12-BM has back-to-back radioactive sample experiments running tomorrow into Thursday afternoon.
  • UNLV will run radioactive samples starting tomorrow at HP-CAT.
  • 20-ID is going to put in a new optics table during the next shutdown.  Wendy is trying to find out who will do floor plate grouting now that Tecknit is gone.  She is going to consult with MOM and possibly Scheck or Central Shops.

Clay White

  • Clay is arranging a meeting with Sector 11 and 12-BM beamline personnel along with Karen Schroeder about beam not being returned to its location at this run’s start up.  After steering was done to correct this issue 11-ID has the best flux they have ever had.  11-BM and 12-BM beamlines have been returned to their original position.  This may be discussed at Ops Directorate.

Bruno Fieramosca

  • The ODH spreadsheet was put into ICMS.
  • Bruno will check 21-ID-G’s ODH monitor sensor today.